### v1

# Parallel Heuristic Exploration for Additive Complexity Reduction in Fast Matrix Multiplication

Abstract
This paper presents a parallel random-search method for reducing additive complexity in fast matrix multiplication. The approach replaces expensive exact evaluation with fast heuristic scoring, including the new Greedy-Intersections strategy. The method runs many independent common subexpression elimination processes in parallel, exploring the search space through random pair substitutions and diverse selection strategies while sharing promising partial solutions. Tested on 164 ternary-coefficient schemes, the method achieves lower addition counts than the state-of-the-art Greedy-Potential on 103 schemes, matches it on 59, and is outperformed on 2. For most schemes, it gives equal or better results while being much faster, making it practical for algorithm exploration. All software and results are open source.
Keywords Fast matrix multiplication $\cdot$ Additive complexity reduction $\cdot$ Common subexpression elimination
1 Introduction
After finding a fast matrix multiplication scheme with low rank, a second important problem appears: reducing the number of additions and subtractions needed to compute it. This additive complexity often determines real-world speed. Common subexpression elimination (CSE) is the standard method for this optimization, but it is a difficult combinatorial problem for the complex expression graphs of fast matrix multiplication algorithms.
Recent work, such as the extensive study by Martensson et al (maartensson2025number), has established benchmarks and introduced methods like “Greedy-Potential” to automate this process. However, the CSE problem for matrix multiplication remains difficult. Greedy strategies often get stuck in local optima, while methods that exactly evaluate each substitution are slow for large schemes.
This paper presents a new approach using heuristic search and massive parallelism, focused on schemes with integer (especially ternary) coefficients. Two implementations are developed: a GPU-accelerated version that runs thousands of processes concurrently for maximum exploration, and a CPU version using OpenMP that runs tens to hundreds of processes for practical use. Both use the same core heuristic strategies. The main contributions are:
  • A family of heuristics for CSE in fast matrix multiplication, with the new Greedy-Intersections strategy as the main component. This heuristic quickly scores candidate substitutions using a stochastic function, balancing immediate gain and future potential without expensive full evaluations.
  • A parallel optimization tool available in both GPU (CUDA) and CPU (OpenMP) versions. The tool runs many independent CSE processes concurrently. Processes use different strategies (rarely including Greedy-Potential as one of them) and share promising partial solutions, allowing broad exploration of the search space.
  • Comprehensive evaluation on 164 known ternary-coefficient ($Z_{T}=\{-1,0,1\}$) schemes. The method achieves lower addition counts than the fmm_add_reduction (fmm_add_reduction2025) baseline on 103 schemes, matches it on 59, and is outperformed on 2. All software and results are open source.
The results show that parallel heuristic search, combining randomness and diverse strategies, can find better reductions than purely greedy or local methods. This advances toward fully automated optimization of matrix multiplication code.
The paper is organized as follows: Section 2 reviews related work. Section 3 describes the methodology. Section 4 presents experimental results. Section 5 discusses limitations and implications. Section 6 concludes. Section 7 describes the availability of software and data.
2 Background and Related Work
2.1 Additive Complexity in Fast Matrix Multiplication
The discovery of a matrix multiplication scheme with low rank is the primary goal for reducing multiplicative complexity. However, to achieve the best possible performance from a given scheme, a secondary optimization is essential: minimizing its additive complexity – the number of required additions and subtractions.
For two schemes with the same rank, the one with fewer additions will execute faster. Therefore, after finding a good low-rank scheme, applying common subexpression elimination (CSE) to minimize its additive cost becomes a critical optimization step for practical implementation. The naive addition count serves as a starting point, while the optimized count, achieved through CSE, represents the real operation cost. For the irregular expression structures in fast matrix multiplication, CSE is a challenging combinatorial optimization problem.
2.2 Prior Work on Additive Minimization
The systematic study of additive minimization has a rich history, particularly for the foundational 2×2×2 case. Strassen’s original algorithm required 18 additions (strassen1969gaussian). Winograd reduced this to 15 (winograd1971multiplication), which was long believed optimal until Karstadt and Schwartz demonstrated that a change of basis could further reduce the count to 12 (karstadt2020matrix). This breakthrough showed that the framework for optimality itself could be expanded, yielding significant practical speedups. The approach of finding new, sparser algebraic representations via basis change, as further developed by Beniamini et al. (beniamini2020sparsifying) and Holtz et al. (holtz2025alternative), constitutes one major strategy for additive reduction.
The second major strategy is common subexpression elimination (CSE), which operates directly on a fixed scheme’s computational graph. The state-of-the-art in automated CSE is the Greedy-Potential algorithm introduced by Martensson et al. (maartensson2025number). This method, implemented in the fmm_add_reduction tool, scores a candidate substitution by its ”potential”: the number of new substitution possibilities created after applying it. This calculation requires a full trial substitution and re-analysis of all two-element subexpressions, which becomes computationally expensive for large schemes.
These strategies represent distinct approaches to the additive complexity problem. Recent work by Stapleton (stapleton202560) demonstrates a different direction: instead of optimizing additions for an existing scheme, a neural network with ternary regularization discovers a new low-rank tensor decomposition from scratch. The resulting scheme is then post-processed with Greedy-Potential CSE, achieving a record of 60 additions for a rank-23 scheme. A similar principle—searching for better algebraic forms before CSE optimization—underlies the flip-graph (kauers2023flip) exploration in previous work, where ternary flips (perminov2025fast) were used to find schemes with minimal naive addition counts before applying CSE.
This paper focuses on improving the CSE strategy itself. While Greedy-Potential is effective, its need for expensive calculations restricts the search scope. The proposed approach instead uses parallel heuristic search. It replaces exact scoring with faster approximate heuristics, including a new Greedy-Intersections strategy. This avoids computational bottlenecks, enabling broader exploration of the search space.
3 Methodology
3.1 Matrix Multiplication Schemes
A matrix multiplication scheme for multiplying matrices $A\in\mathbb{F}^{m\times n}$ and $B\in\mathbb{F}^{n\times p}$ over a field $\mathbb{F}$ with rank $r$ is defined by three coefficient tensors $U\in\mathbb{F}^{r\times m\times n}$, $V\in\mathbb{F}^{r\times n\times p}$, and $W\in\mathbb{F}^{r\times m\times p}$. It computes $r$ intermediate scalar products:
$m_{1}=(u^{(1)}_{11}a_{11}+\dots+u^{(1)}_{mn}a_{mn})\;\cdot\;$ $(v^{(1)}_{11}b_{11}+\dots+v^{(1)}_{np}b_{np})$
$\vdots\;$
$m_{r}=(u^{(r)}_{11}a_{11}+\dots+u^{(r)}_{mn}a_{mn})\;\cdot\;$ $(v^{(r)}_{11}b_{11}+\dots+v^{(r)}_{np}b_{np}),$
and reconstructs the transposed result matrix $C$, where $C=AB$, as:
$c_{ij}=w^{(1)}_{ij}m_{1}+\dots+w^{(r)}_{ij}m_{r}.$
The tensors $(U,V,W)$ must satisfy the Brent equations (brent1970algorithms). A scheme is denoted by its format and rank as $(m,n,p:r)$. This work focuses on schemes where all coefficients in $U,V,W$ are restricted to the set $\{-1,0,1\}$ (ternary coefficients, denoted $Z_{T}$), which are particularly efficient for hardware implementation.
3.2 The Addition Minimization Problem
The additive cost of a scheme originates from computing three independent sets of linear combinations:
  • Input combinations for matrix A ($E_{U}$): $r$ expressions, each combining $m\times n$ variables $a_{ij}$:
    $e^{(l)}_{U}=u^{(l)}_{11}a_{11}+\cdots+u^{(l)}_{mn}a_{mn},\quad\text{for }l=1,\dots,r.$
  • Input combinations for matrix B ($E_{V}$): $r$ expressions, each combining $n\times p$ variables $b_{jk}$.
    $e^{(l)}_{V}=v^{(l)}_{11}b_{11}+\cdots+v^{(l)}_{np}b_{np},\quad\text{for }l=1,\dots,r.$
  • Output reconstruction for matrix C ($E_{W}$): $m\cdot p$ expressions, each combining $r$ variables $m_{l}$ (the intermediate products):
    $e^{(ij)}_{W}=w^{(1)}_{ij}m_{1}+\cdots+w^{(r)}_{ij}m_{r},\quad\text{for }i=1,\dots,m,j=1,\cdots,p.$
The total additive complexity is the sum of the optimized operation counts for these three independent sets:
$\text{Cost}(E_{U})+\text{Cost}(E_{V})+\text{Cost}(E_{W}).$
The optimization of each set is a common, independent task.
3.2.1 Common Subexpression Elimination Task
The optimization of each set $E_{X}$ (where $X\in\{U,V,W\}$) is formulated as a common subexpression elimination problem.
Let $E$ be a set of $n_{e}$ linear expressions in $n_{x}$ input variables $x_{1},x_{2},\dots,x_{n_{x}}$}:
$E=\{e_{1},e_{2},\dots,e_{n_{e}}\},\quad\text{where }e_{i}=\sum_{j=1}^{n_{x}}{\alpha_{ij}\cdot x_{j}},\quad\alpha_{ij}\in\{-1,0,1\}.$
The naive addition count for computing $E$ directly is:
$A_{\text{naive}}(E)=\sum_{i=1}^{n_{e}}\left({\sum_{j=1}^{n_{x}}\left|\alpha_{ij}\right|}-1\right).$
The CSE process iteratively introduces fresh variables $x_{n_{x}+1},x_{n_{x}+2},\dots,x_{n_{x}+n_{f}}$}. Each fresh variable $x_{k}$ is defined as a two-term linear combination of existing variables ($x_{1},\dots,x_{k-1}$):
$x_{k}=x_{p}\pm x_{q}\quad\text{for some }x_{p},x_{q}\in\{x_{1},x_{2},\dots,x_{n_{x}},x_{n_{x}+1},\dots,x_{k-1}\}.$
Upon introducing $x_{k}$, the set of expressions is updated. All occurrences of the subexpression $(x_{p}\pm x_{q})$ within the expressions of $E$ are replaced by the single variable $x_{k}$. This transforms the coefficient matrix: for each expression $e_{i}$, the coefficients for $x_{p}$ and $x_{q}$ are set to zero, and a new coefficient $\alpha_{ik}=\pm 1$ for $x_{k}$ is introduced.
If the pattern $(x_{p}\pm x_{q})$ occurs $c$ times across all expressions, the substitution saves c1c-1 additions: one addition is spent to define $x_{k}$, and the all $c$ occurrences are replaced by a single variable access.
Let $E^{(final)}$ be the expression set after all substitutions, now defined over $n_{x}+n_{f}$ variables. Its naive cost is $A_{\text{naive}}(E^{(final)})$. The total optimized cost is the sum of the additions used to define all fresh variables plus this final naive cost:
$Cost(E)=n_{f}+A_{\text{naive}}\left(E^{(final)}\right).$
The goal of the CSE algorithm is to find a sequence of pair substitutions that gives the smallest total cost $\text{Cost}(E)$.
3.2.2 Illustrative Example
Consider three expressions in variables $x_{1},x_{2},x_{3},x_{4}$}:
$e_{1}$ $=x_{1}+x_{2}-x_{3}+x_{4},$
$e_{2}$ $=x_{1}-x_{2}-x_{4},$
$e_{3}$ $=x_{1}-x_{2}-x_{3}+x_{4}.$
The naive addition count is:
$A_{\text{naive}}=(4-1)+(3-1)+(4-1)=8.$
Introduce fresh variables $x_{5}=x_{2}+x_{4}$ and $x_{6}=x_{1}-x_{3}$. After substitution, the system becomes:
$x_{5}$ $=x_{2}+x_{4},$
$x_{6}$ $=x_{1}-x_{3},$
$e_{1}$ $=x_{5}+x_{6},$
$e_{2}$ $=x_{1}-x_{5},$
$e_{3}$ $=-x_{2}+x_{4}+x_{6}.$
The final expression set $E^{(final)}$ consists of $e_{1},e_{2},e_{3}$} over variables $x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}$}.
Its naive cost is:
$A_{\text{naive}}\left(E^{(final)}\right)=(2-1)+(2-1)+(3-1)=4.$
Applying the cost formula with $n_{f}=2$:
$Cost(E)=n_{f}+A_{\text{naive}}\left(E^{(final)}\right)=2+4=6.$
This represents a reduction from the original 8 to 6 additions, demonstrating a successful application of CSE. The example also shows the updated system, where $x_{5}$ and $x_{6}$ become new variables in the final expressions.
3.3 Core algorithm
The CSE algorithm optimizes each linear expression set $E$ (representing $E_{U}$, $E_{V}$, or $E_{W}$) independently through an iterative process. The procedure for a single set $E$ is described below.
3.3.1 Algorithmic Overview
The core optimization is an iterative loop that terminates when no beneficial substitution can be found. Algorithm 1 shows the complete procedure.
In the Analysis step, pairs are canonized: terms are ordered by index ($i<j$) and the sign is normalized so the coefficient of $x_{i}$ is +1+1. This ensures equivalent forms like $x_{i}-x_{j}$ map to $(x_{i}+x_{j})$. The Selection step uses one of the heuristic strategies described in Section 3.4. The loop repeats until no pair appears more than once.
Input: set of linear expressions $E=\{e_{1},\dots,e_{n_{e}}\}$ in variables $\{x_{1},\dots,x_{n_{x}}\}$
Output: optimized expressions $E^{(final)}$, fresh variables list $F$, total cost
$F\leftarrow\emptyset$
$E_{current}\leftarrow E$
1exrepeat
 Analysis: find all canonical pairs $(x_{i}\pm x_{j})$ in $E_{current}$ and count frequency $c$ for each pair
  1exSelection: choose pair $(x_{i}\pm x_{j})$ with $c>1$ using heuristic strategy
  1exSubstitution:
   Create fresh variable $x_{k}$ with definition $x_{k}=x_{i}\pm x_{j}$
 $F\leftarrow F\cup\{x_{k}\}$
 foreach $e\in E_{current}$ do
      Replace $(x_{i}\pm x_{j})$ with $x_{k}$
      Replace $(-x_{i}\mp x_{j})$ with $x_{k}$
until no pair with frequency >1>1 exists
$E^{(final)}\leftarrow E_{current}$
$\text{Cost}\leftarrow|F|+A_{\text{naive}}(E^{(final)})$
return $E^{(final)},F,\text{Cost}$
Algorithm 1 Common Subexpression Elimination (CSE) algorithm
3.3.2 Algorithmic Scope and Rationale
The algorithm only substitutes pairs of variables. This is enough because any longer common subexpression can be built step by step from pairs. For example, first replace $x_{i}+x_{j}$ with a new variable $x_{t_{1}}$. Then replace $x_{t_{1}}+x_{k}$ with $x_{t_{2}}$. This captures the longer expression $x_{i}+x_{j}+x_{k}$ in two steps.
The main challenge is choosing which pair to substitute next. Always picking the most frequent pair is simple but often leads to poor results. Better selection strategies are needed to explore the search space effectively. The next section describes several strategies, including the new Greedy-Intersections heuristic.
3.4 Heuristic Selection Strategies
The pair selection strategy chooses which common subexpression to eliminate next. Different strategies explore the search space in different ways, often resulting in different final addition counts. This section describes the heuristic strategies used in the parallel search.
3.4.1 Basic Strategies
Four basic strategies provide different balances between exploitation and exploration:
  • Greedy (g): always selects the pair with highest frequency $c$. This maximizes immediate gain but often gets stuck in local optima.
  • Greedy-Alternative (ga): finds all pairs with maximum frequency $c_{\text{max}}$, then selects one uniformly at random. This breaks ties randomly while still exploiting high-frequency pairs.
  • Weighted-Random (wr): selects a pair with probability proportional to its profit c1c-1. More common pairs are more likely to be chosen, but any pair may be selected.
  • Greedy-Random (gr): with probability $p>0.5$, selects a pair using the Greedy-Alternative strategy. Otherwise, selects a pair using the Weighted-Random strategy. This balances exploitation of high-frequency pairs with exploration of other options.
3.4.2 The Greedy-Intersections Heuristic (gi)
The Greedy-Potential method computes the value of a candidate pair by doing a trial substitution. It counts how many new substitution opportunities appear after the change. This evaluation is slow because it must re-analyze all expressions for each candidate pair.
The Greedy-Intersections heuristic estimates the value of a pair $s_{p}=(x_{i}\pm x_{j})$ without doing the substitution. It uses only the current counted pair frequencies. The idea is based on how substitution affects other pairs.
When a pair $s_{p}$ is replaced by a fresh variable $x_{k}$, the effect on another pair $s_{q}$ depends on two factors: whether $s_{p}$ and $s_{q}$ share variables, and whether $s_{q}$ appears in expressions that contain $s_{p}$.
Pairs without intersection ($s_{p}\cap s_{q}=\emptyset$) are unaffected. They keep their current frequency. Pairs with intersection are affected only if they appear in the same expressions as $s_{p}$. If $s_{q}$ appears in expressions that do not contain $s_{p}$, it stays unchanged.
Exact tracking of expression overlap would require full expression analysis. Instead, the heuristic uses a stochastic approximation. For each intersecting pair $s_{q}$, it randomly decides whether $s_{q}$ will survive the substitution. The probability and weight of survival are determined empirically.
The score for pair $s_{p}$ is:
$H(s_{p})=\underbrace{(c_{s_{p}}-1)}_{\text{immediate gain}}+\ \alpha\cdot\sum_{s_{q}\neq s_{p}}I(s_{p},s_{q}),$
where $\alpha$ is a scaling factor chosen randomly from $[0,0.5]$ at the start of each reduction process, and $I(s_{p},s_{q})$ estimates the future value of $s_{q}$:
[…]
Here $\beta$ is a fixed value chosen randomly from $[0.5,1.0]$ at the start of each reduction process. This stochastic approach accounts for the uncertainty in how intersecting pairs evolve after substitution.
The heuristic chooses the pair with maximum $H(s_{p})$. Computing all scores requires checking pairwise intersections, which takes $O(m^{2})$ time for $m$ pairs. This is faster than Greedy-Potential, which needs $O(m^{2}\cdot n_{e})$ time for $n_{e}$ expressions.
3.4.3 Mixed Strategy (mix)
The Mixed strategy selects between available strategies at each step according to predefined weights. The probability of selecting each strategy is proportional to its weight: Greedy-Intersections (weight 8), Greedy-Alternative (weight 4), Greedy-Random (weight 2), Weighted-Random (weight 1). This allows the optimization to adapt to different stages of the search process.
3.5 Parallel Optimization Tool
The CSE problem has many local optima. A single heuristic often finds a good but not optimal solution. To explore the search space more thoroughly, a parallel optimization tool is implemented in both GPU (CUDA) and CPU (OpenMP) versions.
3.5.1 Parallel Reduction Process
For a given matrix multiplication scheme, $N$ independent reduction processes are executed in parallel. In the GPU version, $N$ typically ranges from 2048 to 8192 processes, in the CPU version, from 64 to 256 processes. Each process reduces the three expression sets ($E_{U}$, $E_{V}$, $E_{W}$) independently using the iterative CSE algorithm.
The first process in the first iteration always uses the Greedy strategy as baseline. The remaining processes select their strategy randomly according to configurable weights. The default weights are: Greedy-Intersections (8), Greedy-Alternative (4), Greedy-Random (2), Weighted-Random (1), Mixed (0.1), and Greedy-Potential (0.01). This distribution favors the most promising heuristics while still exploring less common strategies. Users can adjust these weights for specific optimization tasks.
3.5.2 Solution Sharing Mechanism
From the second iteration onward, the tool identifies the best solution found so far (the substitution sequence with the lowest addition count). Then, a configurable percentage (default 40%) of the reduction processes are partially reinitialized. Each reinitialized process copies the first $k$ substitutions from the best solution, where $k$ is chosen randomly between 1 and $3/4$ of the length of the best sequence. The process then continues optimizing from that partial solution.
This allows promising substitution patterns to propagate while maintaining search diversity.
3.5.3 Dynamic Scheme Transformation via Random Flips
As an additional exploration mode, the GPU tool can incorporate transformations from the ternary flip graph (perminov2025fast). This mode was used during initial scheme discovery to find algebraic forms with potential for better additive reduction, though not in the final comparative experiments. The mode operates as follows:
  1. 1.
    Given an input scheme $S$, create $M$ schemes $S_{1},\ldots,S_{M}$ (typically $M=32$).
  2. 2.
    Scheme $S_{1}$ is an exact copy of the original scheme $S$.
  3. 3.
    Each scheme $S_{i}$ for $i=2,\ldots,M$ is created by applying a random number of ternary flip operations to $S$.
  4. 4.
    These $M$ schemes are distributed to the $N$ parallel reduction processes.
  5. 5.
    After the first optimization iteration completes, the process repeats: schemes $S_{2},\ldots,S_{M}$ are regenerated by applying new random flip sequences to the original scheme $S$.
  6. 6.
    The regenerated schemes are redistributed to reduction processes for the next iteration.
This approach continuously explores fresh random transformations of the original scheme throughout the optimization. Each flip operation preserves mathematical correctness while creating different coefficient patterns in the $U$, $V$ and $W$ tensors. Different patterns may offer better opportunities for common subexpression elimination.
The mode explores many alternative schemes with the same rank, searching for forms that are easier to optimize. The best result is a scheme with lower additive complexity while keeping the same rank.
3.6 Implementation Details
The implementation uses a compact representation for linear expressions. An expression $e=\alpha_{1}x_{i_{1}}+\dots+\alpha_{t}x_{i_{t}}$ with coefficients $\alpha_{j}\in\{-1,1\}$ is stored as a set of signed integers. For a variable $x_{i}$:
  • if coefficient +1+1: store +i+i,
  • if coefficient 1-1: store i-i.
Thus expression $x_{1}-x_{4}+x_{5}-x_{9}$ is stored as $\{1,-4,5,-9\}$.
During CSE, substituting a fresh variable $x_{k}$ for the pair $(x_{i}\pm x_{j})$ replaces all occurrences of $(i,\pm j)$ with +k+k and all occurrences of $(-i,\mp j)$ with k-k in the signed integer representation. This requires removing the corresponding signed entries from the set and inserting +k+k or k-k. Set operations (insertion, deletion, membership test) are implemented in $O(1)$ average time using hash tables. Therefore, replacing a pair across all expressions costs $O(n_{e})$ operations where $n_{e}$ is the number of expressions.
This representation is memory-efficient for ternary coefficients and enables fast pair identification and substitution, supporting the parallel execution of thousands of independent CSE processes.
4 Experimental results
4.1 Experimental Setup
The proposed parallel optimization tool is evaluated on 164 fast matrix multiplication schemes with ternary coefficients ($Z_{T}=\{-1,0,1\}$). The schemes include both previously known algorithms and new discoveries obtained through ternary flip graph exploration (perminov2025fast). Starting schemes were selected to have minimal naive addition counts as favorable starting points for additive reduction.
The baseline for comparison is the fmm_add_reduction tool (fmm_add_reduction2025). Three configurations of this baseline are tested:
  • Greedy-Vanilla: simple greedy strategy;
  • Greedy-Potential (5 steps): default greedy-potential mode with 5 values of the $\alpha$ parameter (linearly spaced from 0 to 0.5);
  • Greedy-Potential (40 steps): greedy-potential with 40 values of $\alpha$ (linearly spaced from 0 to 0.5), representing a more exhaustive search.
Two versions of the proposed tool are tested: a GPU version and a CPU version. The GPU version adapts the number of parallel processes to memory constraints:
  • Small schemes (rank $r<100$): 16,384 parallel processes;
  • Medium schemes ($100\leq r<200$): 8,192 parallel processes;
  • Large schemes ($r\geq 200$): 2,048 parallel processes.
The CPU version runs on an Intel(R) Core(TM) i7-8700 CPU @ 3.20GHz and uses fewer processes:
  • Small schemes: 256–512 parallel processes;
  • Medium and large schemes: 32–64 parallel processes.
Each optimization runs until convergence: the best addition count remains unchanged for 10 consecutive iterations. The total number of iterations ranges from 10-15 for small schemes to 50-100 for larger schemes. GPU experiments were performed on an NVIDIA GTX 1650 laptop GPU and an RTX 2050 desktop GPU, both with 6GB memory.
4.2 Additive Complexity Reduction Results
The proposed parallel optimization tool is compared against three baselines: Greedy-Vanilla (GV), Greedy-Potential with 5 $\alpha$ steps (GP-5), and Greedy-Potential with 40 $\alpha$ steps (GP-40). Across 164 ternary-coefficient schemes, the results fall into three categories.
4.2.1 Improved Addition Counts
Table 1 shows schemes where the proposed method achieves lower addition counts than GP-40. These 103 schemes demonstrate the effectiveness of the parallel heuristic search in escaping local optima.
Table 1: Schemes with improved addition counts (proposed method better than GP-40)
Format Rank Naive GV GP GP Prop. Format Rank Naive GV GP GP Prop.
$m\times n\times p$ 5 40 $m\times n\times p$ 5 40
$2\times 3\times 10$ 50 198 135 135 135 134 $4\times 4\times 11$ 134 1150 534 528 524 519
$2\times 3\times 13$ 65 256 170 170 170 169 $4\times 4\times 12$ 145 1495 619 606 596 590
$2\times 4\times 7$ 45 308 181 178 177 174 $4\times 4\times 13$ 157 1474 650 638 635 628
$2\times 4\times 8$ 51 354 197 191 190 188 $4\times 4\times 14$ 169 1596 711 702 694 691
$2\times 4\times 9$ 59 309 210 208 208 207 $4\times 4\times 15$ 181 1679 741 738 735 729
$2\times 4\times 11$ 71 430 275 271 268 265 $4\times 4\times 16$ 192 2059 730 728 724 718
$2\times 4\times 12$ 77 484 282 277 276 274 $4\times 5\times 6$ 90 1023 401 393 386 380
$2\times 4\times 13$ 84 595 323 316 316 313 $4\times 5\times 7$ 104 931 423 413 405 402
$2\times 4\times 14$ 90 616 326 320 318 313 $4\times 5\times 8$ 118 1521 543 538 522 517
$2\times 4\times 15$ 96 662 366 358 357 351 $4\times 5\times 9$ 137 1217 564 555 547 542
$2\times 4\times 16$ 102 708 352 344 342 338 $4\times 5\times 10$ 151 1207 590 580 571 566
$2\times 5\times 5$ 40 283 156 154 154 153 $4\times 5\times 11$ 165 1801 718 702 689 681
$2\times 5\times 6$ 47 332 189 186 184 181 $4\times 5\times 12$ 179 1977 778 772 761 750
$2\times 5\times 7$ 57 340 197 194 192 189 $4\times 6\times 6$ 105 894 467 435 435 430
$2\times 5\times 8$ 65 376 222 218 218 214 $4\times 6\times 7$ 123 1586 535 533 518 517
$2\times 5\times 9$ 72 465 275 270 270 266 $4\times 6\times 8$ 140 1248 576 558 558 551
$2\times 5\times 11$ 87 540 331 326 326 323 $4\times 6\times 9$ 160 1472 720 651 647 646
$2\times 5\times 12$ 94 664 336 330 328 322 $4\times 6\times 10$ 175 1878 759 738 723 715
$2\times 6\times 7$ 68 396 233 230 230 226 $4\times 7\times 7$ 145 1381 680 646 632 629
$2\times 6\times 8$ 77 456 272 269 269 266 $4\times 7\times 8$ 164 1505 752 729 698 690
$2\times 6\times 9$ 86 548 302 296 295 293 $4\times 7\times 9$ 187 2059 806 800 788 785
$2\times 6\times 10$ 94 668 334 327 327 325 $4\times 8\times 8$ 182 1884 861 812 800 795
$2\times 7\times 8$ 90 648 350 333 331 329 $5\times 5\times 5$ 93 843 392 391 387 384
$2\times 7\times 9$ 102 678 398 384 382 379 $5\times 5\times 6$ 110 1215 491 479 466 460
$3\times 3\times 6$ 43 284 168 164 164 160 $5\times 5\times 7$ 127 1607 561 553 539 531
$3\times 3\times 9$ 65 347 218 216 216 215 $5\times 5\times 8$ 144 1924 645 638 631 620
$3\times 3\times 12$ 86 582 303 297 297 289 $5\times 5\times 9$ 167 1814 706 699 690 682
$3\times 3\times 14$ 101 664 373 365 363 361 $5\times 5\times 10$ 184 2116 796 787 781 774
$3\times 4\times 6$ 56 359 211 210 210 209 $5\times 5\times 11$ 202 2272 879 872 859 850
$3\times 4\times 10$ 93 582 351 346 346 345 $5\times 5\times 12$ 220 2444 848 835 811 799
$3\times 4\times 15$ 138 958 544 537 535 533 $5\times 6\times 6$ 130 1716 589 587 573 562
$3\times 5\times 5$ 58 357 230 224 223 221 $5\times 6\times 7$ 150 2039 696 691 671 664
$3\times 5\times 6$ 70 561 270 267 267 265 $5\times 6\times 8$ 170 2312 747 740 728 720
$3\times 5\times 7$ 83 494 315 308 305 303 $5\times 6\times 9$ 197 2376 884 875 850 840
$3\times 5\times 9$ 105 646 390 384 384 383 $5\times 6\times 10$ 217 2772 956 954 940 925
$3\times 5\times 10$ 116 714 402 391 391 389 $5\times 7\times 7$ 176 2605 832 820 799 788
$3\times 5\times 11$ 128 967 497 489 488 484 $5\times 7\times 8$ 206 1880 948 892 882 879
$3\times 5\times 12$ 140 1152 491 483 483 482 $5\times 7\times 9$ 231 2554 993 981 971 963
$3\times 6\times 6$ 85 998 367 364 359 352 $5\times 8\times 8$ 230 2747 1019 1010 980 974
$3\times 6\times 7$ 100 704 364 352 351 350 $6\times 6\times 6$ 153 2182 704 685 671 655
$3\times 6\times 8$ 113 1240 467 459 446 445 $6\times 6\times 7$ 183 2502 810 790 777 769
$3\times 6\times 9$ 127 1095 519 502 493 489 $6\times 6\times 8$ 203 1994 896 880 868 836
$3\times 6\times 10$ 140 1152 494 484 477 476 $6\times 6\times 9$ 225 2440 1029 951 936 923
$3\times 7\times 7$ 115 789 441 434 434 432 $6\times 7\times 7$ 215 2004 965 900 886 880
$3\times 7\times 9$ 147 1002 592 583 580 579 $6\times 7\times 8$ 239 2263 1112 1050 1042 1027
$4\times 4\times 4$ 49 474 167 164 163 159 $6\times 7\times 9$ 268 3062 1184 1166 1152 1146
$4\times 4\times 5$ 61 452 247 237 237 235 $6\times 8\times 8$ 266 2780 1244 1214 1181 1161
$4\times 4\times 6$ 73 540 288 282 281 280 $7\times 7\times 7$ 250 2417 1119 1077 1070 1067
$4\times 4\times 7$ 85 631 326 323 320 319 $7\times 7\times 8$ 279 2926 1369 1260 1248 1237
$4\times 4\times 8$ 96 973 396 391 387 377 $7\times 7\times 9$ 316 3452 1460 1404 1385 1383
$4\times 4\times 9$ 110 925 445 424 422 419 $8\times 8\times 8$ 343 4434 1748 1709 1668 1661
$4\times 4\times 10$ 122 879 493 482 482 480
4.2.2 Equal Performance
Table 2 shows representative schemes where the proposed method matches GP-40. Both methods achieve substantial reductions from naive counts.
Table 2: Schemes with equal addition counts (proposed method matches GP-40)
Format Rank Naive GV GP GP Prop. Format Rank Naive GV GP GP Prop.
$m\times n\times p$ 5 40 $m\times n\times p$ 5 40
$2\times 2\times 2$ 7 24 15 15 15 15 $2\times 4\times 10$ 65 340 223 222 222 222
$2\times 2\times 3$ 11 20 20 20 20 20 $2\times 5\times 10$ 80 416 275 273 273 273
$2\times 2\times 4$ 14 36 31 31 31 31 $2\times 6\times 6$ 57 326 231 228 228 228
$2\times 2\times 5$ 18 38 33 33 33 33 $2\times 7\times 7$ 77 452 323 320 320 320
$2\times 2\times 6$ 21 54 44 44 44 44 $2\times 8\times 8$ 100 608 427 424 424 424
$2\times 2\times 7$ 25 56 46 46 46 46 $3\times 3\times 3$ 23 97 60 60 60 60
$2\times 2\times 8$ 28 72 57 57 57 57 $3\times 3\times 4$ 29 134 93 92 92 92
$2\times 2\times 9$ 32 74 59 59 59 59 $3\times 3\times 5$ 36 193 125 123 123 123
$2\times 2\times 10$ 35 90 70 70 70 70 $3\times 3\times 7$ 51 279 155 154 154 154
$2\times 2\times 11$ 39 92 72 72 72 72 $3\times 3\times 8$ 58 275 170 169 169 169
$2\times 2\times 12$ 42 108 83 83 83 83 $3\times 3\times 10$ 72 386 217 214 214 214
$2\times 2\times 13$ 46 110 85 85 85 85 $3\times 3\times 11$ 79 493 296 287 286 286
$2\times 2\times 14$ 49 126 96 96 96 96 $3\times 3\times 13$ 94 593 341 335 334 334
$2\times 2\times 15$ 53 128 98 98 98 98 $3\times 3\times 15$ 108 579 309 305 305 305
$2\times 2\times 16$ 56 144 109 109 109 109 $3\times 3\times 16$ 115 862 424 410 407 407
$2\times 3\times 3$ 15 58 46 46 46 46 $3\times 4\times 4$ 38 194 136 133 133 133
$2\times 3\times 4$ 20 82 58 58 58 58 $3\times 4\times 5$ 47 277 170 161 161 161
$2\times 3\times 5$ 25 110 72 71 71 71 $3\times 4\times 7$ 64 446 252 251 249 249
$2\times 3\times 6$ 30 116 81 81 81 81 $3\times 4\times 8$ 74 461 272 267 267 267
$2\times 3\times 7$ 35 140 100 99 99 99 $3\times 4\times 9$ 84 535 324 317 316 316
$2\times 3\times 8$ 40 164 104 104 104 104 $3\times 4\times 11$ 102 641 387 384 381 381
$2\times 3\times 9$ 45 174 116 116 116 116 $3\times 4\times 12$ 111 721 424 412 411 411
$2\times 3\times 11$ 55 222 146 145 145 145 $3\times 4\times 13$ 121 804 471 456 455 455
$2\times 3\times 12$ 60 232 151 151 151 151 $3\times 4\times 14$ 128 896 466 460 458 458
$2\times 3\times 14$ 70 280 181 180 180 180 $3\times 4\times 16$ 148 1043 532 518 518 518
$2\times 3\times 15$ 75 307 192 192 192 192 $3\times 5\times 8$ 94 584 295 295 295 295
$2\times 3\times 16$ 80 328 196 196 196 196 $3\times 7\times 8$ 128 930 462 450 450 450
$2\times 4\times 4$ 26 130 93 92 92 92 $3\times 8\times 8$ 148 1020 519 511 511 511
$2\times 4\times 5$ 33 184 114 112 112 112 $4\times 5\times 5$ 76 530 315 300 299 299
$2\times 4\times 6$ 39 202 138 136 136 136
These 59 schemes represent cases where both methods find the same (likely optimal or near-optimal) solution. In most of these cases, GP-5 and GP-40 produce identical addition counts, suggesting these values may indeed be optimal for the given algebraic form. The consistent 35-45% reduction from naive counts demonstrates the effectiveness of CSE optimization.
4.2.3 Inferior Performance
Table 3 shows the 2 schemes where GP-40 achieves lower addition counts than the proposed method. For all of these 2 schemes the proposed method still beats the default GP-5 configuration.
Table 3: Schemes where GP-40 performs better than proposed method
Format Rank Naive GV GP-5 GP-40 Proposed
$6\times 6\times 10$ 252 3540 1291 1249 1167 1210
$7\times 8\times 8$ 310 3604 1670 1498 1471 1494
The performance gap grows with scheme size. The largest differences are 43 additions for $(6,6,10:252)$. However, GP-40 requires running the optimization with 40 different $\alpha$ parameter values, which for large schemes can take tens of hours of computation time. In contrast, the parallel heuristic CPU search completes in 3-4 hours.
These results highlight the combinatorial difficulty of additive complexity reduction. While exhaustive parameter sweeps (GP-40) can find better solutions for very large schemes, they come at significant computational cost. To match GP-40’s results, the proposed method would likely need substantially more computation time. The current implementation provides a favorable trade-off, achieving near-optimal results much faster for most schemes, and still outperforming the practical default configuration (GP-5) in most challenging cases.
4.2.4 Component-Wise Analysis
An analysis of $U$, $V$ and $W$ components separately reveals that different optimization methods can find better reductions in different parts of the same scheme. Table 4 shows schemes where the proposed method and fmm_add_reduction achieve different addition counts for individual components. For these schemes, taking the minimum value from each component across both methods gives a total addition count lower than what either method achieves alone.
Table 4: Component-wise addition counts for schemes improvable by approaches result combination
Format Rank Additions of $U$ Additions of $V$ Additions of $W$ Total additions
$m\times n\times p$ Prop. GP-40 Min Prop. GP-40 Min Prop. GP-40 Min Prop. GP-40 Min
$4\times 7\times 8$ 164 158 160 158 227 234 227 305 304 304 690 698 689
$4\times 8\times 8$ 182 200 205 200 256 258 256 339 337 337 795 800 793
$5\times 7\times 8$ 206 195 194 194 290 290 290 394 398 394 879 882 878
$6\times 6\times 10$ 252 320 289 289 370 356 356 520 522 520 1210 1167 1165
$7\times 7\times 7$ 250 288 288 288 290 289 289 489 493 489 1067 1070 1066
$7\times 8\times 8$ 310 384 387 384 454 444 444 656 640 640 1494 1471 1468
$8\times 8\times 8$ 343 462 465 462 456 465 456 743 738 738 1661 1668 1656
The columns "Min" show the minimum addition count achieved for each component across both methods. This reveals that:
  • The proposed method often finds better reductions for $U$ and $V$ components.
  • fmm_add_reduction sometimes outperforms proposed method in optimizing the $W$ components.
  • Neither method consistently dominates across all three components.
This analysis suggests potential for a hybrid approach: combine the best-optimized U, V and W components from different optimization runs to achieve even lower total addition counts. For example, taking the minimum from each column in Table 4 would reduce totals below what either method achieves alone.
4.2.5 Summary Statistics
Across the 164 ternary-coefficient schemes tested, the proposed parallel method demonstrates consistent improvement over the state-of-the-art:
  • 103 schemes (62.8%): proposed method achieves lower addition counts;
  • 59 schemes (36.0%): proposed method matches GP-40 results;
  • 2 schemes (1.2%): GP-40 achieves better addition counts.
For most cases, the parallel heuristic search provides equal or better results than the exhaustive GP-40 method while being significantly faster. On moderate-sized schemes, the proposed approach typically completes in minutes, while GP-40 can take hours.
For the largest schemes (typically rank >250>250), the pattern changes. The proposed method may require several hours and still yield slightly higher addition counts than GP-40. This highlights the combinatorial difficulty of additive reduction: random exploration works well when the search space is manageable but becomes less effective as dimensionality grows.
Despite this limitation, the proposed method consistently outperforms the practical default configuration GP-5, achieving lower addition counts on 115 schemes (70.1%). This makes it a valuable tool for rapid exploration and optimization of matrix multiplication schemes, especially during algorithm development where fast iteration is important.
5 Discussion and Limitations
The parallel heuristic approach shows that massive exploration with diverse strategies, including randomness and occasional use of expensive methods like Greedy-Potential, can overcome local optima in common subexpression elimination. The success of the Greedy-Intersections heuristic, which avoids expensive trial substitutions, demonstrates that fast approximate scoring can effectively guide the search while maintaining computational efficiency.
5.1 Limitations
Several limitations of the current approach deserve attention:
  • Combinatorial complexity: as scheme size grows, the search space expands exponentially. Although the parallel tool explores thousands of paths, it cannot guarantee optimality. Results show diminishing returns on very large schemes, where exhaustive methods like GP-40 keep an advantage.
  • Heuristic dependence: the quality of results relies on the set of heuristics used. While the current heuristics work well for smaller schemes, larger schemes might benefit from different heuristic designs.
  • Ternary coefficient limitation: the method is developed and evaluated only for schemes with coefficients in $\{-1,0,1\}$. Its effectiveness on schemes with other coefficient sets (such as integer $\mathbb{Z}$ or rational $\mathbb{Q}$) remains untested.
  • Randomness: the stochastic nature means results are not deterministic. Multiple runs can produce different addition counts, though variance is generally low for most schemes.
5.2 Practical Implications
The parallel tool is fast, making it useful for developing matrix multiplication schemes. Researchers can quickly check the addition count of different schemes, allowing them to explore more options. The tool and results are open source, so others can use and improve them.
5.3 Future Work
Several directions for improvement emerge:
  • Adaptive heuristics: strategies that learn from the optimization process could improve performance on large schemes.
  • Hierarchical search: combining the parallel approach with systematic search in promising regions might bridge the gap to exhaustive methods.
  • Integration with basis optimization: joint optimization of basis and CSE could yield further additive reductions.
  • Extension to non-ternary coefficients: generalizing the approach to other coefficient sets would make it useful for more schemes.
6 Conclusion
This paper presents a parallel heuristic tool for reducing additive complexity in fast matrix multiplication. The method addresses the combinatorial challenge of common subexpression elimination by running many optimization processes at once, each using different selection strategies.
The main contributions are:
  • The Greedy-Intersections heuristic, which estimates the value of substitutions more quickly without expensive trial evaluations.
  • A parallel tool in both GPU (CUDA) and CPU (OpenMP) versions that explores the search space using diverse heuristics and shares promising partial solutions.
  • Evaluation on 164 ternary-coefficient schemes showing lower addition counts on 103 schemes, matching on 59, and worse on 2 compared to the state-of-the-art Greedy-Potential method.
The results show that parallel heuristic search, using randomness and multiple strategies, can find better reductions than purely greedy methods while being much faster. The tool offers a good balance between quality and speed, making it useful for algorithm development.
Analysis of $U$, $V$ and $W$ components separately reveals an important insight: different optimization approaches can find better reductions in different parts of the same scheme. Neither method consistently beats the other across all three components, suggesting potential for hybrid approaches that combine the best-optimized components from multiple runs.
For the largest schemes, the problem remains difficult, and exhaustive methods keep a small advantage at higher cost. Still, the proposed tool beats the practical default (GP-5) on 115 out of 164 schemes.
All software and optimized schemes are open source, providing a resource for further work on automating addition reduction in matrix multiplication algorithms. The component-wise analysis data in particular may guide future hybrid optimization strategies.
7 Availability
The complete implementation of the parallel optimization tools, supporting utilities, and all optimized matrix multiplication schemes are publicly available under open-source licenses:
The repositories include documentation, usage examples, and scripts to reproduce the experimental results. The tools can be used as standalone optimizers or integrated into larger scheme exploration workflows.